alones (alones):

3x - 4y = 6 6x - 8y = 10 ;[

OpenStudy (prettygirl_shynice):

3rd problem 3x = 6 + 4y x = (6 +4y) / 3 replace this x in the 2nd equation 6 [ (6+4y) / 3 ] - 8y = 10 2(6+4y) - 8y = 10 12 + 8y -8y = 10

alones (alones):

thansk soo much<3